OpenStudy (awesome781):

How do I simplify: (Square root)7x^2•(Square root)21x^3

OpenStudy (awesome781):

@kmacdaddy

OpenStudy (awesome781):

@phi

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \color{blue} { \sqrt{7x^2} } \times \sqrt{21x^3} }\]correct?

OpenStudy (awesome781):

Yes.

OpenStudy (phi):

can you simplify just the "x" part? what is sqrt(x^2) ?

OpenStudy (solomonzelman):

\[\color{blue}{ \color{blue} { \sqrt{7x^2} } \times \sqrt{21x^3} }=\color{blue}{ \color{blue} { \sqrt{7x^2 \times 21x^3} }}=\color{blue}{ \sqrt{147x^5} }=\color{blue}{ \sqrt{7^3x^5 \times 3} }\]

OpenStudy (solomonzelman):

do it from here.

OpenStudy (solomonzelman):

"So, what's the answer?" you need not the answer, you need to know how to do this.

OpenStudy (solomonzelman):

and that's (not what's my bad) is NOT the answer.

OpenStudy (awesome781):

I know that......I want to know how u got it.....I was trying to know if that was the answerr.